Daisy is running 5 m/s due East while crossing with Joel who is running 4 m/s due North. What is Daisy's actual velocity and angle relative to Joel's?


1
Expert's answer
2021-11-21T17:31:28-0500


The relative velocity of Daisy with respect to Joel


"v_{DJ}=\sqrt{v_D^2+v_J^2}=\sqrt{5^2+4^2}=6.4\:\rm m/s"

The angle

"\theta=\tan ^{-1}\frac{v_J}{v_D}=\tan ^{-1}\frac{5}{4}=51^{\circ}\; \rm S\;of\; E"